Note: Numbers without units are often meaningless. 1. The maximum possible rate for a reaction between two molecules occurs wKs is the dissociation constant
Answer #1

Question 2.

For substrate 1, Ks = k-1/k1 = 103/(7*107) = 1.43*10-5

KM = (k-1 + k2)/k1 = (103 + 2*104)/(7*107) = 3*10-4

Cat. Eff (kcat) = k2 = 2*104 s-1

For substrate 2, Ks = k-1/k1 = 8*105/(3*107) = 0.0267

KM = (k-1 + k2)/k1 = (8*105 + 4*104)/(3*107) = 0.028

Cat. Eff (kcat) = k2 = 4*104 s-1